# FERGI: Automatic Scoring of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### v3

@@ C AU Model Training
This section provides more details of the training of the AU model not provided in Sec. 4.
@@ C.1 Datasets
The DISFA dataset [53] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total. Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5. See Fig. S10 for a visual reference guide for the analyzed AUs.
@@ C.3 Model Training
The weights for the MSE loss and those for the cross entropy loss are both inverse-frequency weighted (for the MSE loss these are based on the merged groups of $\{0,1\}$ and $\{2,3,4,5\}$). Specifically, the weights for the MSE loss are defined as
@@ D Data Filtering
This section provides the details of frame exclusion not provided in Sec. 5.1.

### v4

@@ C AU Model Training
This section provides more details of the training of the AU model not provided in Section 4.
@@ C.1 Datasets
The DISFA dataset [53] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total. Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5. See Figure S11 for a visual reference guide for the analyzed AUs.
@@ C.3 Model Training
The weights for the MSE loss and those for the cross entropy loss are both inverse-frequency weighted and normalized within each AU for addressing class imbalance in the datasets (substantially higher number of occurrences for low AU intensities). However, since the number of occurrences of high intensities are too few for most AUs (resulting in too high weights for the MSE loss if used directly), we “bin” the intensities into 2 groups, and each group shares the same weight. Specifically, for the MSE loss, we apply one weight for occurrences of intensities of 0 and 1 and another weight for occurrences of intensities of 2, 3, 4, and 5, and these weights are computed based on the total number of occurrences within each intensity group.
Specifically, the weights for the MSE loss are defined as
@@ D Data Filtering
This section provides the details of frame exclusion not provided in Section 5.1.
@@
E Experiments
E.1 FAU-Net Analysis
A full version of Figure 5 showing the weights of the hidden nodes in the FAU-Net is shown as Figure S12.
(a) ${\color[rgb]{1,0,0}w_{h(1)-o}\approx-0.58}$
(b) ${\color[rgb]{1,0,0}w_{h(2)-o}\approx-0.53}$
(c) ${\color[rgb]{1,0,0}w_{h(3)-o}\approx-0.50}$
(d) ${\color[rgb]{1,0,0}w_{h(4)-o}\approx-0.49}$
(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
Figure S12: Weights of hidden nodes. Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights). The hidden nodes are ordered by the weights from them to the output node in this figure.